OpenStudy (fanduekisses):

lim x->0 4 * sinxcosx-sinx/x^2

OpenStudy (fanduekisses):

\[\lim_{x \rightarrow 0}\]\[4\frac{ sinxcosx-sinx }{ x^2 }\]

OpenStudy (fanduekisses):

If I just plug in, I get zero :(

OpenStudy (solomonzelman):

you get 0/0 if you plug in 0

OpenStudy (jhannybean):

This is an indeterminant form that requires you to use LH rule.

OpenStudy (fanduekisses):

What is LH rule? I haven't learn that yet.

OpenStudy (jhannybean):

\[4\lim_{s \rightarrow 0}\frac{\sin(x)\cos(x) - \sin(x)}{x^2}\]

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\lim_{x \rightarrow ~0} \frac{\sin x \cos x- \sin x}{x^2}}\) \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~0} \frac{\sin x(\cos x- 1)}{x^2}}\) \(\large\color{black}{\displaystyle\lim_{x \rightarrow ~0} \frac{\sin x}{x} \times \displaystyle\lim_{x \rightarrow ~0} \frac{\cos x- 1}{x}}\)

OpenStudy (jhannybean):

Good :)

OpenStudy (solomonzelman):

so seems as though that it would still be zero:) http://www.wolframalpha.com/input/?i=lim%5Bx-%3E+0%5D+%28sinx+%28cos+x+-+1%29%29%2Fx%5E2

OpenStudy (solomonzelman):

and 0 times 4 is stiill zero:)

OpenStudy (solomonzelman):

I am referring to the rules: ~ \(\large\color{blue}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin(x)}{x}=1}\) ~ \(\large\color{blue}{\displaystyle\lim_{x \rightarrow ~0}\frac{1-\cos(x)}{x}=0}\)

OpenStudy (solomonzelman):

your second limit, in black, is same as the rule number 2 in blue, but except that it is multiplied times -1, but we all know 0 times -1 is still a zero.

OpenStudy (solomonzelman):

@Fanduekisses questions?

OpenStudy (fanduekisses):

Ok so you plugged in, the cos(0) =1 and used the rules

OpenStudy (solomonzelman):

no need plugging anything... I separated the limits, using a rule: `A limit of a product is a product of limits.` then I applied the 2 rules in blue

OpenStudy (fanduekisses):

Where/how did you get the cosx-1 then?

OpenStudy (jhannybean):

they're just trig identities when you take the limit as your variable approaches 0

OpenStudy (jhannybean):

You factor our the \(\sin(x)\).

OpenStudy (solomonzelman):

this site is shi(t) disconnects every time

OpenStudy (solomonzelman):

I will post overview if you want

OpenStudy (fanduekisses):

it was sinxcosx-sinx/x^2 then sinxcosx-1/x^2

OpenStudy (jhannybean):

Not exactly.

OpenStudy (fanduekisses):

oooooooooohhhhhhhh

OpenStudy (fanduekisses):

the sinx/sinx = 1

OpenStudy (jhannybean):

Only when x approaches 0.

OpenStudy (jhannybean):

By The Squeeze Theorem.

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @SolomonZelman I am referring to the rules: ~ \(\large\color{blue}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin(x)}{x}=1}\) ~ \(\large\color{blue}{\displaystyle\lim_{x \rightarrow ~0}\frac{1-\cos(x)}{x}=0}\) \(\color{blue}{\text{End of Quote}}\) Memorize this.

OpenStudy (solomonzelman):

\(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin x\cos x-\sin x}{x^2}}\) factoring the top: \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin x(\cos x-1)}{x^2}}\) re-writing as: \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin x}{x} \times \frac{\cos x-1}{x}}\) limit separation, \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin x}{x} \times \displaystyle\lim_{x \rightarrow ~0}\frac{\cos x-1}{x}}\) \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin x}{x} \times \displaystyle\lim_{x \rightarrow ~0}\frac{-(1-\cos x)}{x}}\) \(\Large\color{black}{\displaystyle\lim_{x \rightarrow ~0}\frac{\sin x}{x} \times (-\displaystyle\lim_{x \rightarrow ~0}\frac{1-\cos x}{x})}\)

OpenStudy (solomonzelman):

is all I did

OpenStudy (solomonzelman):

then apply the blue rules

OpenStudy (solomonzelman):

makes sense?

OpenStudy (jhannybean):

\(0 \cdot -1 = 0\) Typo.

OpenStudy (solomonzelman):

-0 is still a zero 0... is more like it. but it's all same

OpenStudy (jhannybean):

Technically with limits, it means yuo are approaching 0 from the left.

OpenStudy (solomonzelman):

your notation si wrong

OpenStudy (solomonzelman):

don't write two math signs like this: separate with parenthesis:P \(\large\color{black}{ 0 \cdot(-1)=0 }\) like this.

OpenStudy (solomonzelman):

anyways, that ain't matters

OpenStudy (solomonzelman):

the poster went offline, so apparently she got the needed assistance with the question. (wow, I didn't disconnect just now, so unusual !!)